Compute the left and right Riemann sums - L4 and R4, respictively - for

F(x) =(4-x2)1/2 on [-2, 2] and compare their values.
Answer #1

Page And given that f(x) = 2 + 5xt (-4x²) 7(x) = f(2+52-4x²) du f(x) = 5-8 x +15) = 5-885 If15) = -35 Ane Given that Given